OpenStudy (anonymous):

A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total $47,687. The variable costs will be $11.25 per book. The publisher will sell the finished product to bookstores at a price of $22 per book. How many books must the publisher produce and sell so that the production costs will equal the money from sales?

OpenStudy (anonymous):

@mathslover @ganeshie8

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

HINT: Write what you know. Let x = number of books sold Let C(x) = cost of book production Let R(x) = revenue earned by books sold Given: The one-time fixed costs will amount to $29,070. The variable costs will be $9.75 per book. Means: C(x) = 29070 + 9.75x Given: The publisher will sell the finished product to bookstores at a price of $21 per book. Means: R(x) = 21x Given: How many books must the publisher print and sell so that the production costs will equal the money obtained from sales? Means: When does C(x) = R(x) Equation to solve. C(x) = R(x) ===========> 29070 + 9.75x = 21x Subtract 9.75x from both sides. 29070 + 9.75x - 9.75x = 21x - 9.75x 29070 = 11.25x Divide both sides by 11.25. 29070 / 11.25 = 11.25x / 11.25 2584 = x ANSWER: 2584 books must be printed and sold so that the production costs equal the money obtained from sales.

OpenStudy (anonymous):

That's not right.